OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. 1st triangle is 4 & 8 2nd triangle is x & 10

OpenStudy (anonymous):

This is trivial stuff, stop trying to get easy answers when people are trying to seriously help you understand this...

OpenStudy (anonymous):

is it 3.2?

Directrix (directrix):

4/x = 8/10 8x = 40 x = 5

OpenStudy (anonymous):

@Directrix: How do you know that the correspondence is in that order?

OpenStudy (anonymous):

I will say that this question is ill-formed or poorly worded.